OpenStudy (anonymous):

Hello to whom ever can help me, I have an algebra 2 math test tomorrow and was wondering if someone can help me through it. One question is for x and y intercepts: y+4=-2(x-3)

TheSmartOne (thesmartone):

Hello Welcome to Open Study. The x-inercept is when y=o and the y-intercept is when x=0

TheSmartOne (thesmartone):

y-intercept is y+4=-2(0-3) y+4=-2(-3) y+4=6 y=2 So the first point is (2,0)

OpenStudy (anonymous):

OMG stop you're perfect!!! thank you

TheSmartOne (thesmartone):

the x-intercept is 0+4=-2(x-3) 4=-2x+6 -2x=-2 x=1 So the second point would be (1,0)

OpenStudy (anonymous):

Ok thank you very much.